# Take Fake as Real: Realistic-like Robust Black-box Adversarial Attack to Evade AIGC Detection

### v1

@@ IV-G Ablation Experiments
Randomly generated post-processing intensity. We use randomly generated post-processing intensity (without optimization, so the query is N/A) to create adversarial examples for validating the effectiveness of R2BA optimization. The results show that without guidance from the detector’s fake probability, these examples struggle to cross decision boundaries and perform poorly in anti-detection. If the randomly generated intensity is weaker, the adversariality is also weaker, while stronger intensity will lead to severe distortion, prompting the detector to classify the image as AI-generated as well.
R2BA with single post-processing. Due to the different vulnerabilities of detectors to various post-processing techniques, there are performance differences between R2BA with different post-processing. For example, the ASR of R2BA with Gaussian noise is 0.249 on UnivFD(DRCT)[4], but nearly 3 times higher on ConvB(DRCT)[4]. In addition, the anti-detection performance of R2BA with JPEG compression significantly better than that of R2BA with light spot, likely because detectors are more sensitive to high-frequency information (which JPEG compression causes to be lost) than to changes in statistical features (which light spot alters).
@@
Considering all metrics, and comparing with ablated adversarial attack methods, R2BA with 4 post-processing fusion attacks achieves the best results in both attack success rate and query count, while the image quality metrics BRISQUE[42] and SSIM[35] are either the best or close to the best.
@@ IV-H Attacks on Commercial AIGC Detectors
We use Alibaba’s commercial AIGC detection API[43] to compare the effectiveness of the attacks under real-world conditions. For budget and fairness, we set the maximum query to 100, the number of particles in R2BA to 10, and check every 20 queries to determine if the attack succeeds. For white-box attacks, we test adversarial examples generated by attacking the UnivFD[10] and UnivFD(DRCT)[4] detectors, the two most robust detectors across three datasets.
@@ V Conclusion
In this paper, we explore the robustness of GAN-based and diffusion-based AIGC detectors under adversarial example attacks. Given the uncertainty of detectors in real-world scenarios and the discovery that detectors vary in vulnerability to different types of post-processing, we propose a realistic-like robust black-box adversarial attack method, R2BA, based on post-processing fusion optimization. Our method generates adversarial examples that are both robust and maintain high image quality by optimizing post-processing intensity, effectively evading both GAN-based and diffusion-based detection. We conduct extensive experiments on three GAN-based and diffusion-based AIGC datasets, as well as four GAN-based AIGC detectors, three diffusion-based AIGC detectors and one commercial AIGC detection API. The comprehensive experimental results demonstrate that the proposed attack method achieves promising anti-performance in both GAN-based and diffusion-based settings, under the original scenario and the robust scenario, respectively.

### v2

@@ IV-G Ablation Experiments
Randomly generated post-processing intensity. We use randomly generated post-processing intensity (without optimization, so the query is N/A) to create adversarial examples for validating the effectiveness of R2BA optimization. The results show that without guidance from the detector’s fake probability, these examples struggle to cross decision boundaries and perform poorly in anti-detection. If the randomly generated intensity is weaker, the adversariality is also weaker, while stronger intensity will lead to severe distortion, causing the detector to classify the image as AI-generated as well.
R2BA with single post-processing. Due to the different vulnerabilities of detectors to various post-processing techniques, there are performance differences between R2BA with different post-processing. For example, the ASR of R2BA with Gaussian noise is 0.249 on UnivFD(DRCT)[4], but nearly 3 times higher on ConvB(DRCT)[4]. In addition, the anti-detection performance of R2BA with JPEG compression is significantly better than that of R2BA with light spot, likely because detectors are more sensitive to high-frequency information (which JPEG compression causes to be lost) than to changes in statistical features (which light spot alters).
@@
Considering all metrics and comparing with ablated adversarial attack methods, R2BA with 4 post-processing fusion attacks achieves the best results in both attack success rate and query count, while the image quality metrics BRISQUE[43] and SSIM[36] are either the best or close to the best.
@@ IV-H Attacks on Commercial AIGC Detectors
We use Alibaba’s commercial AIGC detection API[44] to compare the effectiveness of the attacks under real-world conditions. For budget and fairness, we set the maximum query to 100, the number of particles in R2BA to 10, and check every 20 queries to determine whether the attack succeeds. For white-box attacks, we test adversarial examples generated by attacking the UnivFD[10] and UnivFD(DRCT)[4] detectors, the two most robust detectors across three datasets.
@@ V Conclusion
In this paper, we explore the robustness of GAN-based and diffusion-based AIGC detectors under adversarial example attacks. Given the uncertainty of detectors in real-world scenarios and the discovery that detectors vary in vulnerability to different types of post-processing, we propose a realistic-like robust black-box adversarial attack method, R2BA, based on post-processing fusion optimization. Our method generates adversarial examples that are both robust and maintain high image quality by optimizing post-processing intensity, effectively evading both GAN-based and diffusion-based detection. We conduct extensive experiments on three GAN-based and diffusion-based AIGC datasets, as well as four GAN-based AIGC detectors, three diffusion-based AIGC detectors and one commercial AIGC detection API. The comprehensive experimental results demonstrate that the proposed attack method achieves promising anti-detection performance in both GAN-based and diffusion-based settings, under the original, robust, and real-world scenarios, respectively.